OpenStudy (sparklestaraa):

Two ocean beaches are being affected by erosion. The table shows the width, in feet, of each beach at high tide measured where 1995 is represented by year 0.

OpenStudy (sparklestaraa):

@DanJS

OpenStudy (danjs):

ok

OpenStudy (danjs):

Both of the erosion patterns are linear, they both change at a constant rate.

OpenStudy (danjs):

I decided that by testing what the slope was for each one over different time intervals.

OpenStudy (danjs):

You want to do the same thing here as the last problem. Create a line equation using two points from each, (x,y) = (year, width)

OpenStudy (danjs):

Then set them equal, and solve for X, the time when both beaches have the same width.

OpenStudy (danjs):

X will be the number of years after 1995, since year zero is given as the year 1995

OpenStudy (danjs):

yeah, you know how to solve it then?

OpenStudy (danjs):

The year will be between time x=11 and x=12 from looking at the table. Which corresponds to 2006 and 2007

OpenStudy (sparklestaraa):

@misty1212 hi can u help finish this?

OpenStudy (sparklestaraa):

@jim_thompson5910 Can u help finish this?

OpenStudy (sparklestaraa):

im not fully understanding this? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

one sec

OpenStudy (sparklestaraa):

alright :)

jimthompson5910 (jim_thompson5910):

the year of 0 corresponds to 100 in the first column so we have the point (0,100) the next row shows us we have the point (5, 90)

jimthompson5910 (jim_thompson5910):

what is the slope of the line through (0,100) and (5, 90) ?

OpenStudy (sparklestaraa):

do i divide?

jimthompson5910 (jim_thompson5910):

at some point, you will

jimthompson5910 (jim_thompson5910):

you first subtract the corresponding coordinates in the same order

OpenStudy (sparklestaraa):

100/-85?

jimthompson5910 (jim_thompson5910):

not quite

OpenStudy (sparklestaraa):

no -100/-85

jimthompson5910 (jim_thompson5910):

subtract the y coords: 100 - 90 = 10 subtract the x coords: 0 - 5 = -5 notice how I'm subtracting in the same order

jimthompson5910 (jim_thompson5910):

divide the differences (ydiff/xdiff) = 10/(-5) = -2

OpenStudy (sparklestaraa):

:D ok makes some sense now

OpenStudy (sparklestaraa):

so that would desrvibe the patterns when u write something that has to do with a certain part you are explaining can u write with part (a,b,or c) that it applies

jimthompson5910 (jim_thompson5910):

well notice how we have 100 decrease to 90 from year 0 to year 5 so it's decreasing by 2 per year you can see this clearly when we go from 80 to 78 (year 10 to year 11)

jimthompson5910 (jim_thompson5910):

that slope of -2 tells you the rate of change

OpenStudy (sparklestaraa):

ok can u wait while i read through what you wrote @jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (sparklestaraa):

oh so the patterns are the average rate of change?

jimthompson5910 (jim_thompson5910):

yeah the pattern tells you how much it's increasing or decreasing (so does the avg rate of change)

OpenStudy (sparklestaraa):

so B is what that we are looking for?

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (sparklestaraa):

Between which years will the beaches have approximately the same width? its between year 11 and 12 right @jim_thompson5910

OpenStudy (sparklestaraa):

sorry 10 to 11 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you had 11 and 12 correct somewhere between those years, the two beaches will have the same width

jimthompson5910 (jim_thompson5910):

it might help to graph the two functions and see where they cross

OpenStudy (sparklestaraa):

and this for c? Assuming these rates remain constant, what can you do to get a better approximation of when the two beaches will have the same width?

jimthompson5910 (jim_thompson5910):

if you find the actual functions for each column, you can use algebra to find the exact point in time when the widths are equal

OpenStudy (sparklestaraa):

so heres all what i wrote tell me if its right

OpenStudy (sparklestaraa):

3. point 1- (x,y)= (0,100) point 2- (x,y)= (5,90) 100-90=10 0-5=-5 divide difference (y diff/ x diff) = 10/(-5) = -2 Decreasing by 2 per year. -2 is the average rate of change. B. Between year 11 and 12 the beaches will have approximately the same width. C. If you find the actual functions for each column, you can use the algebraic techniques to find an exact point in time when widths are equal. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yeah something like that that looks good

OpenStudy (sparklestaraa):

:) thanks so so so much :D

OpenStudy (sparklestaraa):

Can u help with oneeee more then its over :D @jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok